# Train for the Worst, Plan for the Best: Understanding Token Ordering in Masked Diffusions

### v2

@@ 3.1 Benign distributions with hard masking problems
We now describe a simple model of data under which we explore the computational complexity of masking problems.
@@
A latents-and-observations (L&O) distribution is a data distribution $p_{\rm data}$ over sequence of length $L$ with alphabet size $m$ (precisely, $p_{\rm data}$ is over $\{0,\ldots,m\}^{L}$) is specified by a permutation $\pi$ over indices $\{1,2,\ldots,L\}$, number of latent tokens $N$, number of observation tokens $P$ such that $N+P=L$, prior distribution $p_{\textrm{prior}}$ of latent variables over $\{1,\ldots,m\}$ and efficiently learnable observation functions $\mathcal{O}_{1},\ldots,\mathcal{O}_{P}:\{1,\ldots,m\}^{N}\to\Delta(\{0,\ldots,m\})$, [^1] Here efficiently learnable is in the standard PAC sense: given polynomially many examples of the form $(z,y)$ where $z\sim\pi^{n}$ and $y\sim\mathcal{O}_{j}(z)$, there is an efficient algorithm that can w.h.p. learn to approximate $\mathcal{O}_{j}$ in expectation over $\pi^{n}$.
@@
L&O distributions contain two types of tokens: (1) latent tokens and (2) bservation tokens. Intuitively, latent tokens are tokens in the sequence, indexed by $\pi(1),\pi(2),\ldots,\pi(N)$ that serve as “seeds” that provide randomness in the sequence; the remaining tokens, called observation tokens (indexed by $\pi(N+1),\pi(N+2),\ldots,\pi(N+P)$), are determined as (possibly randomized) functions of the latent tokens via $\mathcal{O}_{1},\ldots,\mathcal{O}_{P}$.
Note that by design, order-aware training, e.g. by permuting the sequence so that $\pi$ becomes the identity permutation and then performing autoregressive training, is computationally tractable: predicting $x^{\pi(i)}$ given $x^{\pi(1)},\ldots,x^{\pi(i-1)}$ is trivial when $i\leq N$ as the tokens are independent, and computationally tractable when $i>N$ because $x^{\pi(i)}$ only depends on $x^{\pi(1)},\ldots,x^{\pi(N)}$ and is efficiently learnable by assumption. In contrast, below we will show examples where if one performs order-agnostic training à la MDMs, one will run into hard masking problems with high probability.
First note that if the observations $(\mathcal{O}_{1},\ldots,\mathcal{O}_{P})$ are given by a cryptographic hash function, then the masking problem of predicting $(x^{\pi(1)},\ldots,x^{\pi(L)})$ given $(x^{\pi(N+1)},\ldots,x^{\pi(N+P)})$ is computationally intractable by design because it requires inverting the hash function. While this is a well-known folklore observation regarding the role of token ordering in language modeling, it is not entirely satisfying because this construction is worst-case in nature – in real-world data, one rarely trains on sequences given by cryptographic hash functions. Furthermore, it only establishes hardness for a specific masking pattern which need not be encountered in the course of running the reverse process.
@@
In all these examples, the hardness results we establish hold even if the algorithm knows all of the parameters of $p_{\rm data}$ as well as the observation functions $\mathcal{O}_{1},\ldots,\mathcal{O}_{P}$.
Due to space constraints, here we focus on the following example, deferring two others to Apps. B.1 and B.2.
@@
The complete proof of the proposition is given in Section B.3 but here we provide an overview of techniques used to prove the above result.
Proof overview.
Additional examples of the hardness.
The above setup can also be generalized to capture Bayesian constraint satisfaction problems (Montanari, 2008; Liu et al., 2022), one notable example of which is the stochastic block model (Decelle et al., 2011). There are analogous predictions for the onset of hardness of inference, which can likewise be translated to hardness of masking problems for seemingly benign L&O distributions.
In Appendix B.1 and B.2, we give two more examples of L&O distributions for which order-aware training is tractable yet order-agnostic training of the MDM is computationally hard. First, we consider L&O distributions whose observations are sparse, noisy parities in the latents and deduce hardness for order-agnostic training from the Sparse Learning Parity with Noise assumption (Alekhnovich, 2003). We then consider L&O distributions whose observations are generalized linear models in the latents, and deduce hardness for a large class of efficient algorithms from existing results on Lipschitz hardness (Alaoui & Gamarnik, 2024) for the symmetric binary perceptron (Aubin et al., 2019).
@@ 3.2 Empirical evidence of hardness via likelihoods
Recent studies (Nie et al., 2024; Zheng et al., 2024) have shown that masked diffusion models (MDMs) underperform compared to autoregressive models (ARMs) on natural text data. In this section, we provide evidence that this performance gap is primarily due to the order-agnostic training of MDMs. Since natural text follows a left-to-right token order, we demonstrate that as training deviates from this order, model performance gradually deteriorates.
@@
In other words, the $\pi$-learner predicts the token at position $\pi(i)$ given the clean tokens $x_{0}^{\pi(0)},\ldots,x_{0}^{\pi(i-1)}$ and masked tokens $x_{0}^{\pi(i)},\ldots,x_{0}^{\pi(L-1)}$. If $\pi$ is the identity permutation, this reduces to the standard (left-to-right) autoregressive model. Note that the MDM loss encodes a $\pi$-learner for every permutation $\pi$ because the MDM loss (1) is equivalent to the average loss of those $\pi$-learners over $\pi$ sampled from $\mathrm{Unif}(\mathbb{S}_{L})$:
$\mathcal{L}_{\theta}$ $=-\mathbb{E}_{\pi,x_{0}\sim p_{\rm{data}}}\left[\sum_{i=0}^{L-1}\log p_{\theta}\left(x_{0}^{\pi(i)}\Big{|}x_{0}[\pi\{i,\ldots,L-1\}]\right)\right]$
where $\mathbb{S}_{L}$ denotes the set of all permutations over $\{0,1,\ldots,L-1\}$ (The proof of the above equivalence is given in Appendix E). Therefore, by measuring the ‘hardness’ of each $\pi$-learner, we can probe differences in hardness between arbitrary masking problems and left-to-right masking problems.
@@
In our experiments, the sequence length $L$ is approximately $10^{3}$, so repeating the above for each $\pi$ is infeasible. Instead, we sample $\pi\sim\mathrm{Unif}(\mathbb{S}_{L})$ and examine the scaling law of the $\pi$-learner’s likelihood. We leverage the codebase from (Nie et al., 2024), where the baseline scaling laws of MDM and ARM were introduced. Moreover, given that RoPE has an inductive bias towards left-to-right ordering, we employ a learnable positional embedding layer for all experiments to correct this. Consequently, we also re-run the baseline results, where RoPE was employed. To investigate how the distance between $\pi$ and the identity permutation affects the scaling law, we sample $\pi$ from other distributions interpolating between $\mathrm{Unif}(\mathbb{S}_{L})$ and the point mass at the identical permutation. Further experimental details are provided in Appendix C.1.
@@ 3.3 Error is imbalanced across masking problems
Consider an L&O distribution with $\pi$ given by the identity permutation and where each observation $\mathcal{O}_{j}$ is deterministically given by $\mathrm{NAE}(x_{i_{1}},x_{i_{2}},x_{i_{3}})\triangleq 1-\mathbf{1}[x_{i_{1}}=x_{i_{2}}=x_{i_{3}}]$ for some randomly chosen (prefixed) triples $(i_{1},i_{2},i_{3})\in[N]$.
For an MDM trained on this distribution, we measure the error it achieves on each task $\log p_{\theta}(x_{0}|x_{0}[M])$ via $\mathbb{E}_{x_{0}}\Bigl{\|}\log p_{\theta}(x_{0}|x_{0}[M])-\log p_{\rm data}(x_{0}|x_{0}[M])\Bigr{\|}^{2}$, where $p_{\rm data}(x_{0}|x_{0}[M])$ denotes the Bayes-optimal predictor. Technically, we do not have access to this, so instead we train another MDM for a much larger number of iterations and use this as a proxy. Fig. 2 reveals that prediction tasks for latent positions (light region) exhibit larger errors compared to those for observation positions (dark region).
@@
Here we revisit the text experiment from Section 3.2. Since we do not have access to the Bayes-optimal predictor, we use the metric $\mathbb{E}_{x_{0}\sim p_{\rm{data}}}\left[\sum_{i=0}^{L-1}\log p_{\theta}\left(x_{0}^{\pi(i)}\Big{|}x_{0}[\pi\{i,\ldots,L-1\}]\right)\right]$.
This captures the accumulation of error across subproblems $p_{\theta}\left(x_{0}^{\pi(i)}\Big{|}x_{0}[\pi\{i,\ldots,L-1\}]\right)$, since $p_{\theta}(x_{0}|x_{0}[M])=p_{\rm{data}}(x_{0}|x_{0}[M])$ minimizes this metric. Fig. 2 shows a clear gap between different subproblems.
@@ 4 MDMs can plan around hard problems
Table 1: L&O-NAE-SAT. Adaptive MDM inference achieves better likelihood matching than vanilla MDM inference. Note that naive guessing leads to $75\%$ accuracy, indicating that vanilla inference performs similar or worse than naive guessing.
@@
Figure 4: Generative Perplexity. We compare the resulting generative perplexity (GenPPL) of adaptive vs. vanilla MDM inference. We employ a pretrained $170$M MDM and LLaMA-$7$B (Touvron et al., 2023) as inference and evaluation, respectively. Adaptive MDM inference ( Blue) leads to a substantial reduction in generative perplexity, while maintaining the entropy.
[…]
  • (a)
    Sample a set of masked tokens $\mathcal{S}=\mathcal{F}\left(\theta,x_{t}\right)\subseteq\{i\mid x_{t}^{i}=0\}$.
  • (b)
    For each $i\in\mathcal{S}$, sample $x_{s}^{i}\sim p_{\theta}(x^{i}|x_{t})$.
4.1 Effective design of ordering oracle
We introduce two different oracles, Top-$K$ and Top-$K$ probability margin. Intuitively, both strategies are based on the idea that $S$ should be selected based on how “certain” the model is about each position. We caution that these strategies should not be confused with notions like nucleus sampling in ARMs (Holtzman et al., 2019); the oracles we describe are for selecting the position of the next token to decode, rather than the value, and thus are only meaningful in the context of MDMs.
@@ 4.1 Effective design of ordering oracle
Top-$K$ probability (Zheng et al., 2023). Suppose we want to unmask $K$ positions at time step $t$, i.e., select $|S|=K$. In the Top-$K$ strategy, the uncertainty of a position is estimated by the maximum probability assigned to any value in the vocabulary. More precisely, the certainty at position $i$ is $\max_{j\in\{0,\ldots,m-1\}}p_{\theta}(x^{i}=j|x_{t})$ and $\mathcal{F}(\theta,x_{t})=\text{Top }K\left(\max p_{\theta}(x^{i}|x_{t})\right)$.
@@
Top-$K$ strategy is a good proxy for many tasks and works well in practice (Zheng et al., 2023; Ye et al., 2024; Wang et al., 2024). However, this approach can often provide misleading estimates of uncertainty. Consider when an MDM is confused between two token values, thus assigning them almost equal but high probabilities. In this case, Top-$K$ strategy may still choose to unmask this position, despite its uncertainty. To mitigate this issue, we propose the following alternative strategy.
@@
Top-$K$ probability margin. In this strategy, the uncertainty of a position is instead estimated using the absolute difference between the two most probable values at position $i$. More precisely, if $j_{1}$ and $j_{2}$ are the two most probable values in vocabulary according to $p_{\theta}(x^{i}|x_{t})$ in position $i$, the certainty in the position is given by $|p_{\theta}(x^{i}=j_{1}|x_{t})-p_{\theta}(x^{i}=j_{2}|x_{t})|$ and $\mathcal{F}(\theta,x_{t})=\text{Top }K\left(|p_{\theta}(x^{i}=j_{1}|x_{t})-p_{\theta}(x^{i}=j_{2}|x_{t})|\right)$. When multiple values have similar probabilities at a position, Top-$K$ probability margin will provide a better estimate of the uncertainty of a position, and when there is a single best choice of value then Top-$K$ and Top-$K$ probability margin work similarly.
4.2 Adaptive MDM inference
L&O-NAE-SAT and text data. For the L&O-NAE-SAT distribution defined in Section 3.3, we evaluate the effectiveness of adaptive inference by measuring the accuracy in predicting the observation tokens. Table 1 reveals a clear improvement over vanilla inference. For the text dataset, we evaluate using the standard metric of generative perplexity, by which likelihood is measured by a large language model. We also compute the entropy of the generated samples to ensure both inference strategies exhibit similar levels of diversity. As shown in Fig. 4, we observe a substantial decrease in generative perplexity using adaptive inference. We defer further experimental details to Appendix D.1.
Logic puzzles. We consider two different types of logic puzzles: Sudoku and Zebra (Einstein) puzzles. Intuitively, for Sudoku, some empty (masked) cells are significantly easier to predict than others and we want to choose the cells that are easier to predict during the inference. We evaluate the effectiveness of adaptive MDM inference over vanilla MDM inference in selecting such cells.
To measure the performance of an inference method, we use the percentage of correctly solved puzzles. For both puzzles, we use train and test datasets from (Shah et al., 2024). For the Sudoku puzzle (Table 2) we observe that adaptive MDM inference, in particular Top-$K$ probability margin, obtains substantially higher accuracy (89.49%) compared to vanilla MDM inference (6.88%). Additionally, Top-$K$ probability margin obtains higher accuracy (89.49%) than Top-$K$ (18.51%). As mentioned in Section 4.1, this is because Top-$K$ probability margin more reliably estimates uncertainty when multiple competing values are close in probability at a given position, as is often the case in Sudoku. For the Zebra puzzle, as shown in Table 3, we observe a consistent result: Top-$K$ (98.5%) and Top-$K$ probability margin (98.3%) outperform vanilla MDM inference (76.9%).
@@ 4.2 Adaptive MDM inference
MDM (Top-$K$ probability) 18.51%
MDM (Top-$K$ prob. margin) 89.49%

### v3

@@ 3.1 Benign distributions with hard masking problems
We now describe a simple model of data under which we explore the computational complexity of masking problems and show the contrast between masking problems encountered by MDMs and ARMs.
@@
A latents-and-observations (L&O) distribution is a data distribution $p_{\rm data}$ over sequence of length $L$ with alphabet size $m$ (precisely, $p_{\rm data}$ is over $\{0,\ldots,m\}^{L}$) is specified by a permutation $\pi$ over indices $\{1,2,\ldots,L\}$, number of latent tokens $N$, number of observation tokens $P$ such that $N+P=L$, prior distribution $p_{\textrm{prior}}$ of latent variables over $\{1,\ldots,m\}$ and efficiently learnable observation functions $\mathcal{O}_{1},\ldots,\mathcal{O}_{P}:\{1,\ldots,m\}^{N}\to\Delta(\{0,\ldots,m\})$, [^1] Here efficiently learnable is in the standard PAC sense: given polynomially many examples of the form $(z,y)$ where $z\sim p_{\textrm{prior}}^{N}$ and $y\sim\mathcal{O}_{j}(z)$, there is an efficient algorithm that can w.h.p. learn to approximate $\mathcal{O}_{j}$ in expectation over $p_{\textrm{prior}}^{N}$.
@@
L&O distributions contain two types of tokens: (1) latent tokens and (2) observation tokens. Intuitively, latent tokens are tokens in the sequence, indexed by $\pi(1),\pi(2),\ldots,\pi(N)$ that serve as “seeds” that provide randomness in the sequence; the remaining tokens, called observation tokens (indexed by $\pi(N+1),\pi(N+2),\ldots,\pi(N+P)$), are determined as (possibly randomized) functions of the latent tokens via $\mathcal{O}_{1},\ldots,\mathcal{O}_{P}$. Observe that L&O distributions specified by a permutation $\pi$ have a natural generation order by permutation $\pi$.
Order-aware training
Order-aware training, i.e. by permuting the sequence so that $\pi$ becomes the identity permutation and then performing autoregressive training, is computationally tractable: predicting $x^{\pi(i)}$ given $x^{\pi(1)},\ldots,x^{\pi(i-1)}$ is trivial when $i\leq N$ as the tokens are independent, and computationally tractable when $i>N$ because $x^{\pi(i)}$ only depends on $x^{\pi(1)},\ldots,x^{\pi(N)}$ and is efficiently learnable by assumption. In contrast, below we will show examples where if one performs order-agnostic training à la MDMs, one will run into hard masking problems with high probability.
Order-agnostic training
We first note that if the observations $(\mathcal{O}_{1},\ldots,\mathcal{O}_{P})$ are given by a cryptographic hash function, then the masking problem of predicting $(x^{\pi(1)},\ldots,x^{\pi(L)})$ given $(x^{\pi(N+1)},\ldots,x^{\pi(N+P)})$ is computationally intractable by design because it requires inverting the hash function. While this is a well-known folklore observation regarding the role of token ordering in language modeling, it is not entirely satisfying because this construction is worst-case in nature – in real-world data, one rarely trains on sequences given by cryptographic hash functions. Furthermore, it only establishes hardness for a specific masking pattern which need not be encountered in the course of running the reverse process.
@@
In all these examples, the hardness results we establish hold even if the algorithm knows all of the parameters of $p_{\rm data}$ as well as the observation functions $\mathcal{O}_{1},\ldots,\mathcal{O}_{P}$. Due to space constraints, here we focus on the following example, deferring two others to Apps. B.1 and B.2.
@@
The complete proof of the proposition is given in Section B.4. We also provide a proof outline in Section B.3 for a comprehensive understanding.
@@ 3.2 Empirical evidence of hardness via likelihoods
In the previous section, we provided theoretical evidence that order-aware training is tractable when data has a natural order but the order-agnostic training is not. In this section, we provide empirical evidence to support this claim, using natural text data. Additionally, recent studies (Nie et al., 2024; Zheng et al., 2024) have shown that masked diffusion models (MDMs) underperform compared to autoregressive models (ARMs) on natural text data. In this section, we provide evidence that this performance gap is primarily due to the order-agnostic training of MDMs. Natural text inherently follows a left-to-right token order, and we show that as training deviates from this order, model performance progressively declines.
@@
In other words, the $\pi$-learner predicts the token at position $\pi(i)$ given the clean tokens $x_{0}^{\pi(0)},\ldots,x_{0}^{\pi(i-1)}$ and masked tokens $x_{0}^{\pi(i)},\ldots,x_{0}^{\pi(L-1)}$. If $\pi$ is the identity permutation, this reduces to the standard (left-to-right) autoregressive training. Note that the MDM loss encodes a $\pi$-learner for every permutation $\pi$ because the MDM loss (1) is equivalent to the average loss of those $\pi$-learners over $\pi$ sampled from $\mathrm{Unif}(\mathbb{S}_{L})$:
$\mathcal{L}_{\theta}=-\mathop{\mathbb{E}}_{\begin{subarray}{c}x_{0}\sim p_{\textrm{data}}\\ \pi\sim\textrm{Unif}(\mathbb{S}_{L})\end{subarray}}\left[\sum_{i=0}^{L-1}\log p_{\theta}\left(x_{0}^{\pi(i)}\Big{|}x_{0}[\pi\{i,\ldots,L-1\}]\right)\right],$
where $\mathbb{S}_{L}$ denotes the set of all permutations over $\{0,1,\ldots,L-1\}$. The proof of the above equivalence is given in Appendix E. Therefore, by measuring the ‘hardness’ of each $\pi$-learner, we can probe differences in hardness between arbitrary masking problems and left-to-right masking problems.
@@
In our experiments, the sequence length $L$ is $2048$, so repeating the scaling laws for each $\pi$ is infeasible. Instead, we sample $\pi\sim\mathrm{Unif}(\mathbb{S}_{L})$ and examine the scaling law of the $\pi$-learner’s likelihood. We leverage the codebase from (Nie et al., 2024), where the baseline scaling laws of MDM and ARM were introduced. Moreover, given that RoPE has an inductive bias towards left-to-right ordering, we employ a learnable positional embedding layer for all experiments to correct this. Consequently, we also re-run the baseline results, where RoPE was employed. To investigate how the distance between $\pi$ and the identity permutation affects the scaling law, we consider two interpolating distributions over permutations between $\mathrm{Unif}(\mathbb{S}_{L})$ (i.e, MDM training) and the point mass at the identical permutation (i.e, ARM training). We sample three permutations from the interpolating distribution and $\mathrm{Unif}(\mathbb{S}_{L})$ and plot the scaling law for each of the permutation. Due to space constraints, we provide further experimental details in Appendix C.1.
@@ 3.3 Error is imbalanced across masking problems
Consider an L&O distribution with $\pi$ given by the identity permutation and where each observation $\mathcal{O}_{j}$ is deterministically given by $\mathrm{NAE}(x_{i_{1}},x_{i_{2}},x_{i_{3}})\triangleq 1-\mathbf{1}[x_{i_{1}}=x_{i_{2}}=x_{i_{3}}]$ for some randomly chosen (prefixed) triples $(i_{1},i_{2},i_{3})\in[N]$. For an MDM trained on this distribution, we measure the error it achieves on each task $\log p_{\theta}(x_{0}|x_{0}[M])$ via $\mathbb{E}_{x_{0}}\Bigl{\|}\log p_{\theta}(x_{0}|x_{0}[M])-\log p_{\rm data}(x_{0}|x_{0}[M])\Bigr{\|}^{2}$, where $p_{\rm data}(x_{0}|x_{0}[M])$ denotes the Bayes-optimal predictor. Technically, we do not have access to this, so instead we train another MDM for a much larger number of iterations and use this as a proxy. Fig. 2 reveals that prediction tasks for latent positions (light region) exhibit larger errors compared to those for observation positions (dark region).
@@
Here we revisit the text experiment from Section 3.2. Since we do not have access to the Bayes-optimal predictor, we use the metric $\mathbb{E}_{x_{0}\sim p_{\rm{data}}}\left[\sum_{i=0}^{L-1}\log p_{\theta}\left(x_{0}^{\pi(i)}\Big{|}x_{0}[\pi\{i,\ldots,L-1\}]\right)\right]$. This captures the accumulation of error across subproblems $p_{\theta}\left(x_{0}^{\pi(i)}\Big{|}x_{0}[\pi\{i,\ldots,L-1\}]\right)$, since $p_{\theta}(x_{0}|x_{0}[M])=p_{\rm{data}}(x_{0}|x_{0}[M])$ minimizes this metric. Fig. 2 shows a clear gap between different subproblems.
@@ 4 MDMs can plan around hard problems
Figure 3: Generative Perplexity. We compare the resulting generative perplexity (GenPPL) of adaptive vs. vanilla MDM inference. We employ a pretrained 170170M MDM and LLaMA-77B (Touvron et al., 2023) as inference and evaluation, respectively. Adaptive MDM inference ( Blue) leads to a substantial reduction in generative perplexity, while maintaining the entropy.
(a) Sample a set of masked tokens $\mathcal{S}=\mathcal{F}\left(\theta,x_{t}\right)\subseteq\{i\mid x_{t}^{i}=0\}$. (b) For each $i\in\mathcal{S}$, sample $x_{s}^{i}\sim p_{\theta}(x^{i}|x_{t})$.
4.1 Effective design of ordering oracle
We introduce two different oracles, Top probability and Top probability margin. Intuitively, both strategies are based on the idea that $S$ should be selected based on how “certain” the model is about each position. We caution that these strategies should not be confused with notions like nucleus sampling in ARMs (Holtzman et al., 2019); the oracles we describe are for selecting the position of the next token to decode, rather than the value, and thus are only meaningful in the context of MDMs.
Table 1: L&O-NAE-SAT. Adaptive MDM inference achieves better likelihood matching than vanilla MDM inference. Note that naive guessing leads to $75\%$ accuracy, indicating that vanilla inference performs similarly or worse than naive guessing.
@@ 4.1 Effective design of ordering oracle
Top probability (Zheng et al., 2023).
@@
Suppose we want to unmask $K$ positions at time step $t$, i.e., select $|S|=K$. In the top probability, the uncertainty of a position is estimated by the maximum probability assigned to any value in the vocabulary. More precisely, the certainty at position $i$ is $\max_{j\in\{0,\ldots,m-1\}}p_{\theta}(x^{i}=j|x_{t})$ and $\mathcal{F}(\theta,x_{t})=\text{Top }K\left(\max p_{\theta}(x^{i}|x_{t})\right)$.
@@
Top probability strategy is a good proxy for many tasks and works well in practice (Zheng et al., 2023; Ye et al., 2024; Wang et al., 2024). However, this approach can often provide misleading estimates of uncertainty. Consider when an MDM is confused between two token values, thus assigning them almost equal but high probabilities. In this case, unmasking according to top probability may still choose to unmask this position, despite its uncertainty. To mitigate this issue, we propose the following alternative strategy.
Top probability margin.
@@
In this strategy, the uncertainty of a position is instead estimated using the absolute difference between the two most probable values at position $i$. More precisely, if $j_{1}$ and $j_{2}$ are the two most probable values in vocabulary according to $p_{\theta}(x^{i}|x_{t})$ in position $i$, the certainty in the position is given by $|p_{\theta}(x^{i}=j_{1}|x_{t})-p_{\theta}(x^{i}=j_{2}|x_{t})|$ and $\mathcal{F}(\theta,x_{t})=\text{Top }K\left(|p_{\theta}(x^{i}=j_{1}|x_{t})-p_{\theta}(x^{i}=j_{2}|x_{t})|\right)$. When multiple values have similar probabilities at a position, top probability margin strategy will provide a better estimate of the uncertainty of a position, and when there is a single best choice of value then top probability and top probability margin work similarly.
@@
MDM (Top probability) 18.51%
MDM (Top prob. margin) 89.49%
4.2 Adaptive MDM inference
L&O-NAE-SAT and text data. For the L&O-NAE-SAT distribution defined in Section 3.3, we evaluate the effectiveness of adaptive inference by measuring the accuracy in predicting the observation tokens. Table 1 in the appendix reveals a clear improvement over vanilla inference. For the text dataset, we evaluate using the standard metric of generative perplexity, by which likelihood is measured by a large language model. We also compute the entropy of the generated samples to ensure both inference strategies exhibit similar levels of diversity. As shown in Fig. 3, we observe a substantial decrease in generative perplexity using adaptive inference. We defer further experimental details to Appendix D.1.
Logic puzzles. We consider two different types of logic puzzles: Sudoku and Zebra (Einstein) puzzles. Intuitively, for Sudoku, some empty (masked) cells are significantly easier to predict than others and we want to choose the cells that are easier to predict during the inference. We evaluate the effectiveness of adaptive MDM inference over vanilla MDM inference in selecting such cells. [^2] A prior work (Ye et al., 2024) reported that a 66M MDM with Top-$K$ inference achieves 100% accuracy on Sudoku. Given that a 6M MDM with Top-$K$ only achieves 18.51% on our dataset (Table 2), this suggests that the Sudoku dataset in (Ye et al., 2024) is significantly easier than ours.
To measure the performance of an inference method, we use the percentage of correctly solved puzzles. For both puzzles, we use train and test datasets from (Shah et al., 2024). For the Sudoku puzzle (Table 2) we observe that adaptive MDM inference, in particular, Top probability margin strategy, obtains substantially higher accuracy (89.49%) compared to vanilla MDM inference (6.88%). Additionally, Top probability margin obtains higher accuracy (89.49%) than Top probability strategy (18.51%). As mentioned in Section 4.1, this is because Top probability margin strategy more reliably estimates uncertainty when multiple competing values are close in probability at a given position, as is often the case in Sudoku. For the Zebra puzzle, as shown in Table 3, we observe a consistent result: Top probability (98.5%) and Top probability margin (98.3%) outperform vanilla MDM inference (76.9%).